A wire stretches from the top of a pole 22 ft high to a stake in the ground which is 10 ft from the foot of the pole. Find the length of the wire.
Answer by jojo14344(1513)

We use pythagorean theorem , c^2 = a^2 + b^2
Applicable only for right triangle where the pole obviously is the vertical line (b)= 22ft, and the distance from the stake to the pole which is the horizontal line (a)= 10ft. And finally, the diagonal line (c) will be the length of wire. So substituting we have,
c^2 = (22)^2 + (10)^2
c^2 = 484 + 100
c = sq rt 584
c= 24.20 ft
thank you,
Jojo